A triangular glass prism has a refractive angle of 60 degree. if the minimum deviation by a ray of light passing through the glass is 42 degree. The refractive index of 


1
Expert's answer
2021-03-23T08:13:51-0400
"n=\frac{\sin\frac{A+\delta_m}{2}}{\sin\frac{A}{2}}\\
n=\frac{\sin\frac{60+42}{2}}{\sin\frac{60}{2}}\\
n=\frac{\sin 51}{\sin 30}=1.55"